Why integral t2(-e-t) with limits t= 0, ∞ becomes zero.


1
Expert's answer
2021-05-07T09:58:10-0400

"\bigstar"

Solution


"\boxed{\int_0^\infin t^2(-e^{-t})dt}"


"-\int_0^\infin t^2(e^{-t})dt"


integrate by parts: 

"\u222b\nf\ng\n\u2032\n=\nf\ng\n\u2212\n\u222b\nf\n\u2032\ng"


"f\n=\nt^\n2\n,\t\ng\n\u2032\n=\ne\n^{\u2212\n\nt}\\\\\n\u2193\n steps \t\t\n\\\\\nf\n\u2032\n=\n2\nt\n,\t\ng\n=\n\u2212\ne\n^{\u2212\nt}"



"=\n\u2212\nt^\n2\ne\n^{\u2212\nt}\n\u2212\n\u222b\n\u2212\n2\nt\ne\n^{\u2212\nt}\nd\nt"


Now solving:

"{\displaystyle\int}-2t\mathrm{e}^{-t}\,\mathrm{d}t"

Apply linearity:

"-2{\displaystyle\int}t\mathrm{e}^{-t}\,\mathrm{d}t"


Now solving:

"{\displaystyle\int}t\mathrm{e}^{-t}\,\mathrm{d}t"

Integrate by parts:


"{\int}\mathtt{f}\mathtt{g}' = \mathtt{f}\mathtt{g} - {\int}\mathtt{f}'\mathtt{g}"


"f\n=\nt\n,\t\ng\n\u2032\n=\ne\n^{\u2212\nt}\\\\\n\u2193\n steps\t\t\n\\\\\nf\n\u2032\n=\n1\n,\t\ng\n=\n\u2212\ne\n^{\u2212\nt}\n:"


"=-t\mathrm{e}^{-t}-{\displaystyle\int}-\mathrm{e}^{-t}\,\mathrm{d}t"

Now solving:

"{\displaystyle\int}-\mathrm{e}^{-t}\,\mathrm{d}t"

"Substitute \space 
u
=
−
t
 
⟶
 \frac{
d
u}
{d
t}
=
−
1
 (steps) 
⟶
 
d
t
=
−
d
u
:"

"={\displaystyle\int}\mathrm{e}^u\,\mathrm{d}u"


on solving


"=\mathrm{e}^{-t}"


Plug in solved integrals:


"-t\mathrm{e}^{-t}-{\displaystyle\int}-\mathrm{e}^{-t}\,\mathrm{d}t"

"=-t\mathrm{e}^{-t}-\mathrm{e}^{-t}"


Plug in solved integrals:

"-2{\displaystyle\int}t\mathrm{e}^{-t}\,\mathrm{d}t"

"=2t\mathrm{e}^{-t}+2\mathrm{e}^{-t}"

Plug in solved integrals:


"-t^2\mathrm{e}^{-t}-{\displaystyle\int}-2t\mathrm{e}^{-t}\,\mathrm{d}t"

"=-t^2\mathrm{e}^{-t}-2t\mathrm{e}^{-t}-2\mathrm{e}^{-t}"

Plug in solved integrals:

"-{\displaystyle\int}t^2\mathrm{e}^{-t}\,\mathrm{d}t"

"=t^2\mathrm{e}^{-t}+2t\mathrm{e}^{-t}+2\mathrm{e}^{-t}"


The problem is solved:

"{\displaystyle\int}-t^2\mathrm{e}^{-t}\,\mathrm{d}t"


"=\left(t^2+2t+2\right)\mathrm{e}^{-t}+C"


After solving integral , we got



"\implies" "\boxed{t^2+2t+2\over e^t}" limit 0 to "\infin"


When we put limits


And solve



"\bigstar"

Limit t"_{\rightarrow \infin}" ["t^2+2t+2\over e^t" ]


Limit is "\infin\over\infin"



To solve this


We use L hospital rule , and



"\bigstar"

We will get the value , answer is

"0"